OpenStudy (anonymous):

a colony of bacteria is growing exponentially according to the function below where t is in hours. how many hours will it take for the population to grow to 200? Round to the nearest and do not include units.

OpenStudy (anonymous):

what is the equation the question is referring to?

OpenStudy (anonymous):

B(t) = (4)(e^0.8t) 200 = (4)(e^0.8t) 50 = e^0.8t 0.8t = ln 50 0.8t = 3.912 t = 5 (nearest integer) therefore it takes 5 hours